OpenStudy (anonymous):

Solve B = –4(c – 2d) for d

OpenStudy (anonymous):

B=-4(C-2D) B=-4C+8D 8D=B+4C D=B/8+C/2

OpenStudy (anonymous):

thank you mann

OpenStudy (anonymous):

Glad to help. :)